# The Efficiency Spectrum of Large Language Models: An Algorithmic Survey

### v1

@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Shanahan, 2022; Huang and Chang, 2022; Chang et al., 2023; Yang et al., 2023a), characterized by their massive scale of tens or even hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2022; Anil et al., 2023), have become a central focus in the field of artificial intelligence. These models, exemplified by applications like ChatGPT (cha, [n. d.]) and Claude (cla, [n. d.]), have demonstrated impressive capabilities in a variety of general-purpose tasks, such as text summarization (Yang et al., 2023b), translation (Hendy et al., 2023), question answering (Robinson and Wingate, 2022), and even rudimentary coding (Chen et al., 2021b), largely attributed to their expertise in natural language understanding. Although the exact mechanisms driving their exceptional performance remain elusive (Zhu and Li, 2023), it is widely believed that their large size endows them with emergent abilities (Wei et al., 2022a) not observed in smaller models, and this is seen as a key step toward achieving Artificial General Intelligence (AGI) (agi, [n. d.]; Bubeck et al., 2023).
@@
While the large size of LLMs is crucial for their capabilities (see Figure 1), it also presents a significant drawback: their deployment is severely limited by high computational costs and memory requirements (Treviso et al., 2023; Zhuang et al., 2023; Zhu et al., 2023b; Xu and McAuley, 2023). The resources needed for training these models are substantial, creating challenges in both resource allocation and model design. For example, the cost of exploring different architectures or strategies becomes prohibitive (Zhao et al., 2023). Furthermore, their large size makes them unsuitable for resource-constrained environments like edge devices, thus narrowing their range of applications (Ahmed and Wahed, 2020). This computational burden also confines the development of LLMs to large companies with abundant resources (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023). Many essential details, such as the data collection pipeline and training methodologies, remain proprietary, which hinders academic research and poses challenges for smaller companies. Additionally, the environmental impact of training these models is not to be overlooked, raising concerns about carbon emissions and ethical considerations (Vinuesa et al., 2020; Wu et al., 2022; Van Wynsberghe, 2021). Consequently, there is a growing emphasis on improving the efficiency of LLMs.
Refer to caption (a) 8-billion model
Refer to caption (b) 62-billion model
Refer to caption (c) 540-billion model
@@
Figure 1. Capability and performance tree (pal, [n. d.]) of PaLM (Chowdhery et al., 2022) across different scales (8-billion, 62-billion, 540-billion). Each circle node represents a specific capability, and its size indicates the corresponding performance level—the larger the circle, the greater the capability. As the model scale increases, performance not only improves across existing tasks but also reveals new capabilities.
@@ 2.1. Core Concepts of LLMs
Architectural Foundations. LLMs can generally be categorized into two main paradigms: encoder-decoder models (Peters et al., 2018; Liu et al., 2019; Sanh et al., 2019; Lee et al., 2020; Lan et al., 2020; Raffel et al., 2020a), exemplified by BERT (Kenton and Toutanova, 2019), and decoder-only models (Radford et al., 2019; OpenAI, 2023; Brown et al., 2020; Radford et al., 2018; Zhang et al., 2022b; Chowdhery et al., 2022; Anil et al., 2023; Scao et al., 2022; Du et al., 2022; Rae et al., 2021; Hoffmann et al., 2022; Thoppilan et al., 2022; Touvron et al., 2023a, b; Wu et al., 2023a), such as the GPT series (OpenAI, 2023; Radford et al., 2018, 2019; Brown et al., 2020). BERT is trained using masked language modeling, enabling it to excel in contextual understanding by predicting masked or missing tokens. On the other hand, GPT models are trained using autoregressive modeling, which equips them with strong generative capabilities by predicting subsequent tokens in a sequence. Despite these differences, both types of models commonly rely on the Transformer (Vaswani et al., 2017) architecture, which is particularly noteworthy for its self-attention mechanism. In self-attention, each token is represented as a key, value, and query. The query weighs the importance of other tokens, represented as keys in understanding a particular token. These weights are applied to the values to create context-aware representations. This mechanism allows each token in the sequence to consider all other tokens simultaneously, facilitating parallel processing of sequential data and effective capture of long-sequence dependencies. As a result, multi-head attention layers are often stacked to form deep networks in LLMs. Nowadays, decoder-only models like GPT-4 (OpenAI, 2023) and LLaMa (Touvron et al., 2023a, b) are becoming increasingly prevalent, yet the core architectural module of self-attention remains a constant across these variations.
@@
Training Essentials. LLMs acquire their general-purpose capabilities from an initial pre-training phase on expansive and diverse datasets (Brown et al., 2020; Yang et al., 2023a). These datasets cover a broad spectrum of sources such as books, scientific papers, code, and websites (Zha et al., 2023). This foundational knowledge is then fine-tuned on relatively smaller datasets in a supervised manner, aimed at enabling the LLMs to adhere to human instructions, a process known as instruction tuning (Christiano et al., 2017; Ziegler et al., 2019; Zhang et al., 2023b; Ouyang et al., 2022; Wang et al., 2022a; Taori et al., 2023; Chiang et al., 2023; Liu et al., 2023a; Zhu et al., 2023a; Dai et al., 2023; Su et al., 2023). A reinforcement learning with human feedback is then proceeded to a instructed fine-tuned LLM to further align model behavior upon human preferences and instructions (Touvron et al., 2023b). In the current landscape, decoder-only models have become the norm, particularly for their superior generative abilities. These models employ autoregressive objectives in the pre-training phase to maximize the likelihood of predicting subsequent tokens based on their preceding context. The scaling (Kaplan et al., 2020; Hoffmann et al., 2022) of this autoregressive pre-training significantly enhances LLM capabilities, as demonstrated by models like the GPT and PaLM series (Anil et al., 2023; Chowdhery et al., 2022). For instance, PaLM (Chowdhery et al., 2022) was trained on a 780B-token dataset and utilized a 540B-parameter Transformer architecture. PaLM-2 (Anil et al., 2023), its successor, advances this further; its largest variant has 1.1T parameters and was trained on a more diverse, multilingual dataset with 1.5T tokens. However, this scaling introduces its own set of challenges, requiring efficient training infrastructures and optimization strategies. Distributed training frameworks that focus on data parallelism, such as DeepSpeed (Rasley et al., 2020) and Fully Sharded Data Parallel (FSDP) (fsd, [n. d.]), or pipeline parallelism like Gpipe (Huang et al., 2019) and PipeDream (Narayanan et al., 2019), are commonly employed. Additionally, tensor parallelism techniques like Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b) and SARATHI (Agrawal et al., 2023) are also utilized. Specialized techniques, such as mixed-precision training (Scao et al., 2022; Nair et al., 2023; Dettmers et al., 2022) and quantization-aware training (Liu et al., 2023b; Xiao et al., 2023; Lin et al., 2023), are often incorporated to streamline the training process. These methods not only address the computational challenges tied to the scale of LLMs but also facilitate the development of increasingly capable models.
@@ 3.1. Introduction
Recent research in predicting large language model (LLM) performance has concentrated on understanding the scaling law. This law delineates how LLM performance is influenced by factors such as model architecture, neural model size, computing power for training, and available data. The concept of scaling law, rooted in statistical mechanics approaches for predicting model generalization, has a rich history dating back to the early 1990s (Györgyi and Tishby, 1990; Amari et al., 1992; Seung et al., 1992; Barkai et al., 1993). Its relevance has been reinvigorated recently in the context of modern deep learning models (Hestness et al., 2017; Rosenfeld et al., 2019; Henighan et al., 2020; Kaplan et al., 2020; Hernandez et al., 2021; Tay et al., 2022a; Muennighoff et al., 2023; Hoffmann et al., 2022; Caballero et al., 2022; Alabdulmohsin et al., 2022; Tay et al., 2021; Sorscher et al., 2022). This section will delve into the latest advancements and insights in the scaling law as applied to LLMs, highlighting how these models evolve and perform under varying conditions.
@@ 5.2. Efficient Attention
The Transformer model, as introduced by Vaswani et al. (Vaswani et al., 2017), utilizes a vanilla attention mechanism that computes dense pairwise relations in the input sequence. This results in quadratic complexity. However, recognizing that not all these relations hold equal significance, recent research has focused on methods to streamline this process. These methods aim to identify and maintain only the most crucial relations, thereby enhancing the efficiency of the attention calculation (Child et al., 2019; Dao et al., 2022b; Dao, 2023; Liu and Abbeel, 2023; Ainslie et al., 2023; Zhuang et al., 2022; Ding et al., 2023b). In this subsection, we include the discussion of two main branches: (i) the use of fast or sparse attentions (Child et al., 2019; Kitaev et al., 2021; Choromanski et al., 2021; Dai et al., 2019; Martins et al., 2022; Ding et al., 2023b; Zhuang et al., 2023; Ma et al., 2023b) and (ii) IO-aware attention calculation with hardware co-design (Dao et al., 2022b; Dao, 2023; Ham et al., 2021, 2020). Both approaches reduce hardware loading time for efficient attention calculations.
@@
Moving away from directly calculating pairwise attention, some methods (Dao et al., 2022a; Fu et al., 2023; Ding et al., 2023b; Liu and Abbeel, 2023) explore the possibility of computing attention at a block level, which enables parallelization of the computation, significantly enhancing efficiency. For instance, the Monarch framework (Dao et al., 2022a) and its advanced version, Monarch Mixer (M2) (Fu et al., 2023), adopt a novel strategy. They sparsify the dense attention matrix by decomposing it into a combination of permutation and block diagonal matrices. This decomposition allows for more efficient processing of attention calculations. Furthermore, the Blockwise Self-Attention (BST) method (Liu and Abbeel, 2023) introduces blockwise computation for both self-attention and feed-forward networks. This technique aims to lower memory requirements typically associated with traditional attention mechanisms. Moreover, some methods like LongNet (Ding et al., 2023b) incorporate dilated attention in place of the original dense attention, allowing the processing of much longer token sequences, thereby expanding the capabilities of LLMs.
Hardware-Related Efficient Attention. Along with designing more efficient attention mechanisms at the software level, a significant focus has shifted to optimizing these mechanisms at a hardware level. One of the main challenges in this domain is efficiently utilizing computational resources, such as High Bandwidth Memory (HBM) and Static Random-Access Memory (SRAM), on GPUs. In this regard, recent advancements like FlashAttention (Dao et al., 2022b) and its successor, FlashAttention-2 (Dao, 2023), have emerged. FlashAttention reconsiders the attention computation by adopting an I/O-centric perspective. It minimizes data transfers between HBM and SRAM, addressing a critical bottleneck in GPU processing. This method integrates block-wise softmax value calculations and updates with statistics. Such integration eliminates the conventional requirement of computing the softmax function only after all attention values have been determined. Building upon FlashAttention (Dao et al., 2022b), FlashAttention-2 (Dao, 2023) further optimizes work partitioning and reduces non-matrix multiplications, capitalizing on the GPUs’ optimization for matrix operations. These algorithms are tailored to hardware considerations and accelerate the models on GPU machines.
@@ 5.5. Attention-free
One significant drawback of the vanilla attention mechanism (Vaswani et al., 2017) is the quadratic complexity of attention computation, making it especially inefficient for handling long sequences. Although efficient / sparse attention offers some relief, its worst-case theoretical complexity remains unchanged. To address this, various attention-free methods have been proposed, providing alternatives that avoid the computation of the quadratic attention matrix (Peng et al., 2023a; Dao et al., 2022c; Poli et al., 2023; Sun et al., 2023a; Zhai et al., 2021; Gu et al., 2021). These methods could be largely categorized into those that replace the attention mechanism with recurrent computation (Zhai et al., 2021; Peng et al., 2023a; Sun et al., 2023a), and those that discretize state space representations (Gupta et al., 2022; Gu et al., 2021, 2022; Mehta et al., 2023). Notably, these new methods like RWKV (Peng et al., 2023a), H3 (Dao et al., 2022c), Hyena (Poli et al., 2023), and RetNet (Sun et al., 2023a) achieve performance comparable to the standard Transformer. RWKV (Peng et al., 2023a) utilizes recurrent neural networks (RNNs) to streamline sequence processing, thereby reducing the complexity of handling long sequences. H3 (Dao et al., 2022c), based on state space models (SSMs), offers an efficient alternative for data representation and processing. Hyena (Poli et al., 2023) presents itself as a drop-in replacement for traditional attention mechanisms, simplifying the Transformer architecture. RetNet (Sun et al., 2023a) introduces a multi-scale retention module coupled with a feed-forward network module, enhancing parallelism and recurrence, which significantly improves efficiency in both training and inference phases. We include the complexity analysis of these methods compared with the vanilla Transformer with an input query of length $n$ in Table 1. It provides an overview of how each method scales in complexity, offering insights into the advancements in the attention-free technologies.
Table 1. Comparisons of parallelization in training time and time/memory cost in inference evaluation between Transformer and attention-free methods when using a sequence of length $n$ as input. Models with more $+$ under performance are of better performance.
Method
Parallelization
Time Cost
Memory Cost
Performance
Transformer (Vaswani et al., 2017)
Yes
$O(n)$
$O(n^{2})$
RWKV (Peng et al., 2023a)
No
$O(1)$
$O(n)$
H3 (Dao et al., 2022c)
Yes
$O(1)$
$O(n\log n)$
Hyena (Poli et al., 2023)
Yes
$O(n)$
$O(n\log n)$
RetNet (Sun et al., 2023a)
Yes
$O(1)$
$O(n)$

### v2

@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Shanahan, 2022; Huang and Chang, 2022; Chang et al., 2023; Yang et al., 2023a), characterized by their massive scale of tens or even hundreds of billions of parameters (Brown et al., 2020; Chowdhery et al., 2022; Anil et al., 2023), have become a central focus in the field of artificial intelligence. These models, exemplified by applications like ChatGPT (OpenAI, 2023b) and Claude (cla, 2023), have demonstrated impressive capabilities in a variety of general-purpose tasks, such as text summarization (Yang et al., 2023b), translation (Hendy et al., 2023), question answering (Robinson and Wingate, 2022), and even rudimentary coding (Chen et al., 2021b), largely attributed to their expertise in natural language understanding. Although the exact mechanisms driving their exceptional performance remain elusive (Zhu and Li, 2023), it is widely believed that their large size endows them with emergent abilities (Wei et al., 2022a) not observed in smaller models, and this is seen as a key step toward achieving Artificial General Intelligence (AGI) (agi, 2023; Bubeck et al., 2023).
@@
While the large size of LLMs is crucial for their capabilities (see Figure 1), it also presents a significant drawback: their deployment is severely limited by high computational costs and memory requirements (Treviso et al., 2023; Zhuang et al., 2023; Zhu et al., 2023b; Xu and McAuley, 2023). The resources needed for training these models are substantial, creating challenges in both resource allocation and model design. For example, the cost of exploring different architectures or strategies becomes prohibitive (Zhao et al., 2023). Furthermore, their large size makes them unsuitable for resource-constrained environments like edge devices, thus narrowing their range of applications (Ahmed and Wahed, 2020). This computational burden also confines the development of LLMs to large companies with abundant resources (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023a). Many essential details, such as the data collection pipeline and training methodologies, remain proprietary, which hinders academic research and poses challenges for smaller companies. Additionally, the environmental impact of training these models is not to be overlooked, raising concerns about carbon emissions and ethical considerations (Vinuesa et al., 2020; Wu et al., 2022; Van Wynsberghe, 2021). Consequently, there is a growing emphasis on improving the efficiency of LLMs.
(a) 8-billion model
(b) 62-billion model
(c) 540-billion model
@@
Figure 1. Capability and performance tree (pal, 2022) of PaLM (Chowdhery et al., 2022) across different scales (8-billion, 62-billion, 540-billion). Each circle node represents a specific capability, and its size indicates the corresponding performance level—the larger the circle, the greater the capability. As the model scale increases, performance not only improves across existing tasks but also reveals new capabilities.
@@ 2.1. Core Concepts of LLMs
Architectural Foundations. LLMs can generally be categorized into two main paradigms: encoder-decoder models (Peters et al., 2018; Liu et al., 2019; Sanh et al., 2019; Lee et al., 2020; Lan et al., 2020; Raffel et al., 2020a), exemplified by BERT (Kenton and Toutanova, 2019), and decoder-only models (Radford et al., 2019; OpenAI, 2023a; Brown et al., 2020; Radford et al., 2018; Zhang et al., 2022b; Chowdhery et al., 2022; Anil et al., 2023; Scao et al., 2022; Du et al., 2022; Rae et al., 2021; Hoffmann et al., 2022; Thoppilan et al., 2022; Touvron et al., 2023a, b; Wu et al., 2023a), such as the GPT series (OpenAI, 2023a; Radford et al., 2018, 2019; Brown et al., 2020). BERT is trained using masked language modeling, enabling it to excel in contextual understanding by predicting masked or missing tokens. On the other hand, GPT models are trained using autoregressive modeling, which equips them with strong generative capabilities by predicting subsequent tokens in a sequence. Despite these differences, both types of models commonly rely on the Transformer (Vaswani et al., 2017) architecture, which is particularly noteworthy for its self-attention mechanism. In self-attention, each token is represented as a key, value, and query. The query weighs the importance of other tokens, represented as keys in understanding a particular token. These weights are applied to the values to create context-aware representations. This mechanism allows each token in the sequence to consider all other tokens simultaneously, facilitating parallel processing of sequential data and effective capture of long-sequence dependencies. As a result, multi-head attention layers are often stacked to form deep networks in LLMs. Nowadays, decoder-only models like GPT-4 (OpenAI, 2023a) and LLaMa (Touvron et al., 2023a, b) are becoming increasingly prevalent, yet the core architectural module of self-attention remains a constant across these variations.
@@
Training Essentials. LLMs acquire their general-purpose capabilities from an initial pre-training phase on expansive and diverse datasets (Brown et al., 2020; Yang et al., 2023a). These datasets cover a broad spectrum of sources such as books, scientific papers, code, and websites (Zha et al., 2023). This foundational knowledge is then fine-tuned on relatively smaller datasets in a supervised manner, aimed at enabling the LLMs to adhere to human instructions, a process known as instruction tuning (Christiano et al., 2017; Ziegler et al., 2019; Zhang et al., 2023b; Ouyang et al., 2022; Wang et al., 2022a; Taori et al., 2023; Chiang et al., 2023; Liu et al., 2023a; Zhu et al., 2023a; Dai et al., 2023; Zhong et al., 2023; Su et al., 2023). A reinforcement learning with human feedback is then proceeded to a instructed fine-tuned LLM to further align model behavior upon human preferences and instructions (Touvron et al., 2023b). In the current landscape, decoder-only models have become the norm, particularly for their superior generative abilities. These models employ autoregressive objectives in the pre-training phase to maximize the likelihood of predicting subsequent tokens based on their preceding context. The scaling (Kaplan et al., 2020; Hoffmann et al., 2022) of this autoregressive pre-training significantly enhances LLM capabilities, as demonstrated by models like the GPT and PaLM series (Anil et al., 2023; Chowdhery et al., 2022). For instance, PaLM (Chowdhery et al., 2022) was trained on a 780B-token dataset and utilized a 540B-parameter Transformer architecture. PaLM-2 (Anil et al., 2023), its successor, advances this further; its largest variant has 1.1T parameters and was trained on a more diverse, multilingual dataset with 1.5T tokens. However, this scaling introduces its own set of challenges, requiring efficient training infrastructures and optimization strategies. Distributed training frameworks that focus on data parallelism, such as DeepSpeed (Rasley et al., 2020) and Fully Sharded Data Parallel (FSDP) (fsd, 2023), or pipeline parallelism like Gpipe (Huang et al., 2019) and PipeDream (Narayanan et al., 2019), are commonly employed. Additionally, tensor parallelism techniques like Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b) and SARATHI (Agrawal et al., 2023) are also utilized. Specialized techniques, such as mixed-precision training (Scao et al., 2022; Nair et al., 2023; Dettmers et al., 2022) and quantization-aware training (Liu et al., 2023b; Xiao et al., 2023; Lin et al., 2023), are often incorporated to streamline the training process. These methods not only address the computational challenges tied to the scale of LLMs but also facilitate the development of increasingly capable models.
@@ 3.1. Introduction
Recent research in predicting large language model (LLM) performance has concentrated on understanding the scaling law (Kaplan et al., 2020). This law delineates how LLM performance is influenced by factors such as model architecture, neural model size, computing power for training, and available data. The concept of scaling law, rooted in statistical mechanics approaches for predicting model generalization, has a rich history dating back to the early 1990s (Györgyi and Tishby, 1990; Amari et al., 1992; Seung et al., 1992; Barkai et al., 1993). Its relevance has been reinvigorated recently in the context of modern deep learning models (Hestness et al., 2017; Rosenfeld et al., 2019; Henighan et al., 2020; Kaplan et al., 2020; Hernandez et al., 2021; Tay et al., 2022a; Muennighoff et al., 2023; Hoffmann et al., 2022; Caballero et al., 2022; Alabdulmohsin et al., 2022; Tay et al., 2021; Sorscher et al., 2022). This section will delve into the latest advancements and insights in the scaling law as applied to LLMs, highlighting how these models evolve and perform under varying conditions.
@@ 5.2. Efficient Attention
The Transformer model, as introduced by Vaswani et al. (Vaswani et al., 2017), utilizes a vanilla attention mechanism that computes dense pairwise relations in the input sequence. This results in quadratic complexity. However, recognizing that not all these relations hold equal significance, recent research has focused on methods to streamline this process. These methods aim to identify and maintain only the most crucial relations, thereby enhancing the efficiency of the attention calculation (Child et al., 2019; Dao et al., 2022b; Dao, 2023; Liu and Abbeel, 2023; Liu et al., 2023d; Ainslie et al., 2023; Zhuang et al., 2022; Ding et al., 2023b). In this subsection, we include the discussion of two main branches: (i) the use of fast or sparse attentions (Child et al., 2019; Kitaev et al., 2021; Choromanski et al., 2021; Dai et al., 2019; Martins et al., 2022; Ding et al., 2023b; Zhuang et al., 2023; Ma et al., 2023b) and (ii) IO-aware attention calculation with hardware co-design (Dao et al., 2022b; Dao, 2023; Ham et al., 2021, 2020; Dao et al., 2023; Hong et al., 2023). Both approaches reduce hardware loading time for efficient attention calculations.
@@
Moving away from directly calculating pairwise attention, some methods (Dao et al., 2022a; Fu et al., 2023; Ding et al., 2023b; Liu and Abbeel, 2023) explore the possibility of computing attention at a block level, which enables parallelization of the computation, significantly enhancing efficiency. For instance, the Monarch framework (Dao et al., 2022a) and its advanced version, Monarch Mixer (M2) (Fu et al., 2023), adopt a novel strategy. They sparsify the dense attention matrix by decomposing it into a combination of permutation and block diagonal matrices. This decomposition allows for more efficient processing of attention calculations. Furthermore, the Blockwise Self-Attention (BST) method (Liu and Abbeel, 2023; Liu et al., 2023d) introduces blockwise computation for both self-attention and feed-forward networks. This technique aims to lower memory requirements typically associated with traditional attention mechanisms. Moreover, some methods like LongNet (Ding et al., 2023b) incorporate dilated attention in place of the original dense attention, allowing the processing of much longer token sequences, thereby expanding the capabilities of LLMs.
Hardware-Related Efficient Attention. Along with designing more efficient attention mechanisms at the software level, a significant focus has shifted to optimizing these mechanisms at a hardware level. One of the main challenges in this domain is efficiently utilizing computational resources, such as High Bandwidth Memory (HBM) and Static Random-Access Memory (SRAM), on GPUs. In this regard, recent advancements like FlashAttention (Dao et al., 2022b) and its successor, FlashAttention-2 (Dao, 2023), have emerged. FlashAttention reconsiders the attention computation by adopting an I/O-centric perspective. It minimizes data transfers between HBM and SRAM, addressing a critical bottleneck in GPU processing. This method integrates block-wise softmax value calculations and updates with statistics. Such integration eliminates the conventional requirement of computing the softmax function only after all attention values have been determined. Building upon FlashAttention (Dao et al., 2022b), FlashAttention-2 (Dao, 2023) further optimizes work partitioning and reduces non-matrix multiplications, capitalizing on the GPUs’ optimization for matrix operations. These algorithms are tailored to hardware considerations and accelerate the models on GPU machines. Based on FlashAttention, FlashDecoding (Dao et al., 2023) and FlashDecoding++ (Hong et al., 2023) split the keys/values in smaller chunks for parallelization of partial attention, and FlashDecoding++ (Hong et al., 2023) further introduce asynchronized softmax with unified max value and flat GEMM optimization with double buffering for further speed up, along with the heuristic dataflow for adaption on hardware resources.
@@ 5.5. Attention-free
One significant drawback of the vanilla attention mechanism (Vaswani et al., 2017) is the quadratic complexity of attention computation, making it especially inefficient for handling long sequences. Although efficient / sparse attention offers some relief, its worst-case theoretical complexity remains unchanged. To address this, various attention-free methods have been proposed, providing alternatives that avoid the computation of the quadratic attention matrix (Peng et al., 2023a; Dao et al., 2022c; Poli et al., 2023; Sun et al., 2023a; Zhai et al., 2021; Gu et al., 2021). These methods could be largely categorized into those that replace the attention mechanism with recurrent computation (Zhai et al., 2021; Peng et al., 2023a; Sun et al., 2023a), and those that discretize state space representations (Gupta et al., 2022; Gu et al., 2021, 2022; Mehta et al., 2023; Gu and Dao, 2023; Smith et al., 2023). Notably, these new methods like RWKV (Peng et al., 2023a), H3 (Dao et al., 2022c), Hyena (Poli et al., 2023), RetNet (Sun et al., 2023a) and Mamba (Gu and Dao, 2023) achieve performance comparable to the standard Transformer. RWKV (Peng et al., 2023a) utilizes recurrent neural networks (RNNs) to streamline sequence processing, thereby reducing the complexity of handling long sequences. H3 (Dao et al., 2022c), based on state space models (SSMs), offers an efficient alternative for data representation and processing. Hyena (Poli et al., 2023) presents itself as a drop-in replacement for traditional attention mechanisms, simplifying the Transformer architecture. RetNet (Sun et al., 2023a) introduces a multi-scale retention module coupled with a feed-forward network module, enhancing parallelism and recurrence, which significantly improves efficiency in both training and inference phases. Mamba (Gu and Dao, 2023) includes the selection operation based on state space models for as compression and further improve the efficiency with hardware-related optimization. We include the complexity analysis of these methods compared with the vanilla Transformer with an input query of length $n$ in Table 1. It provides an overview of how each method scales in complexity, offering insights into the advancements in the attention-free technologies.
Table 1. Comparisons of time and memory cost in inference evaluation between Transformer and attention-free methods when using a sequence of length $n$ as input. Models with more $+$ under performance are of better performance of the perplexity numbers on the in-domain validation set and other out-of-domain corpora.
Method Time Cost Memory Cost Performance
Transformer (Vaswani et al., 2017) $O(n)$ $O(n^{2})$ $+$(+)
RWKV (Peng et al., 2023a) $O(1)$ $O(n)$ +
H3 (Dao et al., 2022c) $O(1)$ $O(n\log n)$ ++
Hyena (Poli et al., 2023) $O(n)$ $O(n\log n)$ +
RetNet (Sun et al., 2023a) $O(1)$ $O(n)$ $+$
Mamba (Gu and Dao, 2023) $O(1)$ $O(n)$ $+$